OpenStudy (anonymous):

If x-y=y-z=z-t=5, what is the value of x-t??

OpenStudy (anonymous):

Where did you get 12?

OpenStudy (mathmate):

It is a telescopic set of equations: x-y=5 y-z=5 z-t=5 Add the three equations together, y & z cancel out to give x-t=15

OpenStudy (anonymous):

Holly hell, what kind of math is this?

OpenStudy (anonymous):

Is this Calculus ?

OpenStudy (anonymous):

On the PSAT's..-____-